HETAV PATEL

Grade 11,

Factorise the following expressions ax^2 y+bxy^2 +cxyz

Factorise the following expressions
ax^2 y+bxy^2 +cxyz

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

ax^2 y+bxy^2 +cxyz
ax^2y = a×x×x×y
bxy^2= b×x×y×y
cxyz = c×x×y×z
Common factors of ax^2y ,bxy^2and cxyz are xy Now, ax^2y+bxy^2+cxyz = xy(ax+by+cz)

Thanks